OpenStudy (anonymous):

Help!!! Find dy/dx : a. y= log(base2) (3x) b. x^2y^4 = ln x + 3y

OpenStudy (anonymous):

For the part "a" can I use Chain Rule to solve ??

OpenStudy (cwrw238):

b. x^2 4y^3 * dy/dx + 2x y^4 = 1/x + 3 * dy/x solve for dy/dx

OpenStudy (cwrw238):

for a. i think you need to the log to base e

OpenStudy (cwrw238):

i'm having a hard time remembering formula for change of base try googling it

OpenStudy (cwrw238):

* convert the log to base e

OpenStudy (anonymous):

Me either.....I don't remember the formula.

OpenStudy (anonymous):

On part b. which term I will divide on both side ???

OpenStudy (loser66):

\[ y' = (\log_2{3x})'\\\frac{dy}{dx}=\frac{3}{xln2}\]

OpenStudy (loser66):

sorry, RHS doesn't have 3 in numerator

OpenStudy (anonymous):

okay.....than will be dy/dx= x ln2 ??

OpenStudy (loser66):

1/xln2

OpenStudy (anonymous):

ok...thank you Loser 66. I appreciate it !!!!

OpenStudy (cwrw238):

in part b isolate all the terms containing dy/dx then factor dy/dx out and then divide

OpenStudy (loser66):

In part b, I think it's implicit differentiate .

OpenStudy (loser66):

like what cwrw238 said.

OpenStudy (anonymous):

I am trying........

OpenStudy (anonymous):

Hello guys.....I still need help on part b.

OpenStudy (loser66):

\[x^2y^4=lnx+3y\\x^2y^4-3y=lnx\] derivative both sides \[x^2*4y^3*y' -3y'=\frac{1}{x}\] factor y' out \[y'(4x^2y^3-3) =\frac{1}{x}\\y'=\frac{1}{x(4x^2y^3-3)}\] or \[\frac{dy}{dx}= \frac{1}{x(4x^2y^3-3)}\]

OpenStudy (anonymous):

Thanks again Loser66.......I tried a lot.....unsuccessfully !!!

OpenStudy (loser66):

yw